OpenStudy (kimberlystanley1617):

Solve the system y = -x + 7 and y = 0.5(x - 3)2

OpenStudy (kimberlystanley1617):

help pleaseeeeee

rishavraj (rishavraj):

use substitution method...u see u got y = -x + 7 so plug tht value of y in 2nd equation

OpenStudy (kimberlystanley1617):

I forgot how to plug it in.... can you please explain how you plug it in.?

rishavraj (rishavraj):

-x + 7 = 0.5(x - 3)2

OpenStudy (kimberlystanley1617):

Thank you